OpenStudy (anonymous):

Convert to Trigonometric Form, Round to hundredth degree. 7-24i (Please show steps to help me understand)

OpenStudy (anonymous):

ok you are going to write this as \[r(cos(\theta)+i\sin(\theta))\]

OpenStudy (anonymous):

you need r and \[\theta\]

OpenStudy (anonymous):

yeah rcis

OpenStudy (anonymous):

r is the easy part.

OpenStudy (anonymous):

\[r=\sqrt{7^2+24^2}\] whatever that is

OpenStudy (anonymous):

25

OpenStudy (anonymous):

use a calculator if you want a decimal

OpenStudy (anonymous):

oh ok then that was really easy

OpenStudy (anonymous):

now to find \[\theta\] use \[\tan(\theta)=\frac{b}{a}=\frac{24}{-7}\]

OpenStudy (anonymous):

-73.73

OpenStudy (anonymous):

put your calculator in degree mode and find \[\tan^{-1}(-\frac{24}{7})\]

OpenStudy (anonymous):

-73.74* my bad

OpenStudy (anonymous):

that wasn't too bad

OpenStudy (anonymous):

i will take your word for it. if the negative angle offends you you can add 360 degrees to it to get a positive number, but your answer is correct

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

so i add 360 to it? cause the book answer is 286.26

OpenStudy (anonymous):

there are an infinite number of ways to write it, so i guess that is what the book did, add 360

OpenStudy (anonymous):

there is only one thing you have to be careful of. you have to know what quadrant you are in

OpenStudy (anonymous):

so i will always use tangent?

OpenStudy (anonymous):

let me be precise.

OpenStudy (anonymous):

it is \[\tan(\theta)=\frac{b}{a}\] always but that doesn't always mean that \[\theta = \tan^{-1}(\frac{b}{a})\]

OpenStudy (anonymous):

because arctan only gives you answers in quadrants 1 and 4, not in 2 or 3

OpenStudy (anonymous):

in this case we were in quadrant 4 because it was 7-24i, 4 over, down 24

OpenStudy (anonymous):

hm what would i have to do for Q2 and 3?

OpenStudy (anonymous):

but if you are in quadrant 2 or 3 you cannot just take the arctangent. you have to adjust to make sure you are in the right quadrant

OpenStudy (anonymous):

well suppose you wanted to do this with \[-7+24i\] everything is still the same

OpenStudy (anonymous):

r is still the same, and arctangent is still the same. but you are in quadrant 2 so 286.26 is obviously wrong

OpenStudy (anonymous):

you would minus 180 right?

OpenStudy (anonymous):

so if you got out your calculator and got -73.74 for your angle, you would have to add 180 degrees

OpenStudy (anonymous):

yes. add or subtract. either way you got it

OpenStudy (anonymous):

106.26

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

ok i believe you. but it is clear yes? because arctan is confined to -90 to 90 so you have to adjust if you are not in quadrant 1 or 4

OpenStudy (anonymous):

hope this helps because the steps are fairly simple

OpenStudy (anonymous):

yeah i understand now thanks a bunch the book example is really weak and doesnt explain well glad i found this site

OpenStudy (anonymous):

yw!